OpenStudy (anonymous):

2sin^2x-cosx-1=0 solution set

OpenStudy (anonymous):

Do you have a guess on the answer or do you need help in explanation.

OpenStudy (dape):

If you replace the 1 in your equation with \(1=\sin^2x+\cos^2x\) and remember the identity \(\cos^2x-\sin^2x=\cos2x\) then you get a simple equation only involving \(\cos x\). Then you only need to take arccos of both sides to get the solution set.